# Yong Ji1, Junye Li1 and Rui Yang∗2,3

### v3

@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
#$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
@@
This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing nonempty closed subsets of $X$. Let
@@
For $p\in\mathbb{F}$ and any $k\geq 1$, we have $p\in\mathbb{F}_{(k,0)}$. Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<3\epsilon_{(k,0)}$. This implies that the orbit of $p$ is $4\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have
@@
For $p\in\mathbb{F}$ and sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
Case $(1)$. There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
@@
$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{c_{k,Q_{k}}}{n}D(\mathcal{E}_{c_{k,Q_{k}}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast})$
$+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast}).$
Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.14), we have $D(\mathcal{E}_{c_{k,Q_{k}}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$. Thus,
@@
Case $(2)$. There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
@@
$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{c_{k,j}}{n}D(\mathcal{E}_{c_{k,j}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
$+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)}).$
@@
Case $(3)$. There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
@@
$+\frac{c_{k-1,Q_{k-1}}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})+\frac{c_{k,0}}{n}D(\mathcal{E}_{c_{k,0}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
@@
$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
@@
and
$D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ $\leq D(\mathcal{E}_{c_{k-1,Q_{k-1}}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})$
$\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
@@
Case $(4)$. There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
@@
$\mathcal{E}_{n}(z)=$ $\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}}{n}\mathcal{E}_{c_{k,j-1}}(f^{M_{(k,j-1)^{\ast}}}(p))$
@@
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on the above arguments, one can also show that the sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ holds for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
@@
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying the $\rho$-measure of some appropriate dynamical balls decays exponentially fast along the sequence $\{M_{k}\}$, where $M_{k}:=M_{(k,Q_{k})}$. Then we can invoke the Entropy Distribution Principle developed in Lemma 2.2 to estimate the lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
@@
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus one has $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$. This implies that $\rho(\mathbb{F})=1$. Recall that $M_{(0,0)}=n_{(0,0)}=1$, by (3.12) and (3.13), we have
#$\#\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
@@
Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be any Bowen ball that intersects $\mathbb{F}$. Then for any $k>k_{0}$, one has
@@
Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not=\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also nonempty by the construction of $\mathbb{F}$ and #$\#(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$. Notice that $\mathbb{F}_{k_{0}+1}$ descends from $\mathbb{F}_{k_{0}}$. Then one has
#$\#(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
@@
For sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
@@ 3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
The following lemma gives an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$, which was already obtained in [Y23, Lemma 3.3,(3)].
@@
Let $(X,f)$ be a t.d.s., and let $K\subset M_{f}(X)$ be a nonempty closed subset.
@@
Thus, by the Lemma 3.6 the upper bound of the Bowen metric mean dimension of saturated set is given by:
@@ 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof for the lower bound is similar to that in subsection 3.1.2. So we only give a sketch for the estimation of the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$.
We show that for any small enough $\epsilon^{\ast}>0$, it holds that
@@
Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination of ergodic $f$-invariant probability measures with rational coefficients such that
@@
Applying Katok’s $\epsilon$-entropy formula and repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
@@
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated and
#$\#E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
@@
#$\#\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
@@
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For $(k,i)\in\mathcal{A}^{\prime}$, and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
@@
Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$. Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every $n$ large enough and any ball $B_{n}(z,\frac{\epsilon^{\ast}}{2})$ intersecting with the fractal set $\mathbb{F}$, it holds that
@@
Then, by the Lemma 2.1, we have
@@
With the help of Theorems 1.1 and 1.2, we prove Theorem 1.3.
Firstly, we show
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\{\underline{\rm mrid}_{\mu}(f,K,d)\}.$ (3.18)
Fix $\mu\in K$ and $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$. Then, by Lemma 3.6, for every $\epsilon>0$,

### v4

@@ 3.1.2 Lower bound of $\overline{\rm mdim}^{P}_{M}(f,G_{K}^{\omega,U},d)$
$\#\mathcal{I}_{(k,j)}=\#\mathcal{I}_{(k,j)^{\ast}}\cdot\#\mathcal{S}_{(k,j)}=\Pi_{(k^{\prime},j^{\prime})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},j^{\prime})}=\Pi_{(k^{\prime},Q_{k^{\prime}})\leq(k,j)}\#\mathcal{S}_{(k^{\prime},Q_{k^{\prime}})}.$ (3.13)
@@
This implies that $\{\mathbb{F}_{(k,j)}:(k,j)\in\mathcal{A}\}$ is a sequence of decreasing non-empty closed subsets of $X$. Let
@@
For $p\in\mathbb{F}$, we have $p\in\mathbb{F}_{(k,0)}$ for any $k\geq 1$. Hence, for any $a_{i}^{k}\in\Delta_{k}$, $1\leq i\leq q_{k}$, there exists $t\in[M_{(k,0)^{\ast}},M_{(k,0)}]$ such that $d(f^{t}(p),a_{i}^{k})<4\epsilon_{(k,0)}$. This implies that the orbit of $p$ is $5\epsilon_{(k,0)}$-dense in $X$. By the arbitrariness of $k\geq 1$ and $p\in\mathbb{F}$, we have
@@
Let $p\in\mathbb{F}$. For sufficiently large $n\in\mathbb{N}$, to estimate $D(\mathcal{E}_{n}(p),K)$ we split into four cases to discuss.
Case $(1)$: There exists $k>1$ such that $M_{(k+1,0)}\leq n<M_{(k+1,1)}^{1}$.
@@
$D(\mathcal{E}_{n}(p),\mu^{\ast})\leq$ $\frac{M_{(k,Q_{k})^{\ast}}}{n}D(\mathcal{E}_{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})+\frac{n-M_{(k,Q_{k})}}{n}D(\mathcal{E}_{n-M_{(k,Q_{k})}}(f^{M_{(k,Q_{k})}}(p)),\mu^{\ast})$
$+\frac{c_{k,Q_{k}}+m_{(k,Q_{k})}}{n}D(\mathcal{E}_{c_{k,Q_{k}}+m_{(k,Q_{k})}}(f^{M_{(k,Q_{k})^{\ast}}}(p)),\mu^{\ast}).$
Noticing that both $\frac{M_{(k,Q_{k})^{\ast}}}{n}$ and $\frac{n-M_{(k,Q_{k})}}{n}$ are smaller than $\frac{1}{k}$, and using formula (3.14), we have $D(\mathcal{E}_{c_{k,Q_{k}}+m_{(k,Q_{k})}}f^{M_{(k,Q_{k})^{\ast}}}(p),\mu^{\ast})<\frac{7}{k}+2\delta_{(k,Q_{k})}$. Thus,
@@
Case $(2)$: There exists $(k,j)\in\mathcal{A}$ with $j\neq 0$ such that $M_{(k,j)}\leq n<M_{(k,j)_{\ast}}^{1}$.
@@
$D(\mathcal{E}_{n}(p),\mu_{(k,j)})\leq$ $\frac{M_{(k,j)^{\ast}}}{n}D(\mathcal{E}_{M_{(k,j)^{\ast}}}(p),\mu_{(k,j)})+\frac{n-M_{(k,j)}}{n}D(\mathcal{E}_{n-M_{(k,j)}}(f^{M_{(k,j)}}(p)),\mu_{(k,j)})$
$+\frac{c_{k,j}+m_{(k,j)}}{n}D(\mathcal{E}_{c_{k,j}+m_{(k,j)}}(f^{M_{(k,j)^{\ast}}}(p)),\mu_{(k,j)})$
@@
Case $(3)$: There exist $(k,1)\in\mathcal{A}$ and $i>1$ such that $M_{(k,1)}^{i}\leq n<M_{(k,1)}^{i+1}$.
@@
$+\frac{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}{n}D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$
$+\frac{c_{k,0}+m_{(k,0)}}{n}D(\mathcal{E}_{c_{k,0}+m_{(k,0)}}(f^{M_{(k-1,Q_{k-1})}}(p)),\mu_{(k,1)})$
@@
$\frac{M_{(k-1,Q_{k-1})^{\ast}}}{n}<\frac{1}{k-1},\ \frac{c_{k,0}+m_{(k,0)}}{n}<\frac{1}{k},\ \frac{n-M_{(k,1)}^{i}}{n}<\frac{1}{k},$
@@
and $D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu_{(k,1)})$ is smaller than
$D(\mathcal{E}_{c_{k-1,Q_{k-1}}+m_{(k-1,Q_{k-1})}}(f^{M_{(k-1,Q_{k-1})^{\ast}}}(p)),\mu^{\ast})+D(\mu^{\ast},\mu_{(k,1)})\leq\frac{8}{k-1}+2\delta_{(k-1,Q_{k-1})}.$
@@
Case $(4)$: There exist $(k,j)\in\mathcal{A}$ with $j>1$ and $i>1$ such that $M_{(k,j)}^{i}\leq n<M_{(k,j)}^{i+1}$.
@@
$\mathcal{E}_{n}(z)$ $=\frac{M_{(k,j-1)^{\ast}}}{n}\mathcal{E}_{M_{(k,j-1)^{\ast}}}(p)+\frac{c_{k,j-1}+m_{(k,j-1)}}{n}\mathcal{E}_{c_{k,j-1}+m_{(k,j-1)}}(f^{M_{(k,j-1)^{\ast}}}(p))$
@@
To sum up, whatever case we consider, by letting $n\to\infty$ we get $V_{f}(p)\subset K$. Based on the above arguments, one can also show that the sequences $\{\mathcal{E}_{n}(p)\}_{n\in\mathbb{N}}$ and $\{\mu_{(k,j)}\}_{(k,j)\in\mathcal{A}}$ have the same limit points. Notice that $K=\overline{\{\mu_{(k^{\prime},j^{\prime})}:(k^{\prime},j^{\prime})>(k,j)\}}$ for any fixed $(k,j)\in\mathcal{A}$. Consequently, we have $V_{f}(p)=K$, and hence $\mathbb{F}\subset G_{K}$. ∎
@@
Now we define a probability measure $\rho$ on $\mathbb{F}$ satisfying the $\rho$-measure of some appropriate dynamical balls decays exponentially fast along the sequence $\{M_{k}\}$, where $M_{k}:=M_{(k,Q_{k})}$. Then we can invoke the Entropy Distribution Principle stated in Lemma 2.2 to obtain a lower bound of $\epsilon$-packing topological entropy of $\mathbb{F}$.
@@
For simplicity, we write $\mathcal{I}_{k}$, $\mathbb{F}_{k}$ instead of $\mathcal{I}_{(k,Q_{k})}$, $\mathbb{F}_{(k,Q_{k})}$ respectively. Since $M(X)$ is compact, we may assume that $\rho_{k}\to\rho\in M(X)$ as $k\to\infty$. It is clear that $\rho_{k}(\mathbb{F}_{k})=1$ for every $k$, and for any $k>k_{0}$, one has $\rho_{k}(\mathbb{F}_{k_{0}})=1$ since $\mathbb{F}_{k}\subset\mathbb{F}_{k_{0}}$. Thus, $\rho(\mathbb{F}_{k_{0}})\geq\limsup_{k\to\infty}\rho_{k}(\mathbb{F}_{k_{0}})=1$. This implies that $\rho(\mathbb{F})=1$. Recall that $M_{(0,0)}=n_{(0,0)}=1$. By (3.12) and (3.13), we have
$\mathcal{I}_{k}\geq\exp\left[(M_{k}-1)\cdot(1-\zeta)^{2}\cdot(H-3\zeta)\right].$
@@
Let $k_{0}\geq 1$, $z\in X$ and $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ be a Bowen ball that intersects $\mathbb{F}$. Then, for any $k>k_{0}$, one has
@@
Since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}\not$$\emptyset$, the set $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}_{k_{0}}$ is also non-empty by the construction of $\mathbb{F}$ and $(\mathbb{F}_{k_{0}}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))=1$. Notice that $\mathbb{F}_{k_{0}+1}$ descends from $\mathbb{F}_{k_{0}}$. Then one has
$(\mathbb{F}_{k_{0}+1}\cap B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2}))\leq\#\mathcal{S}_{(k_{0}+1,Q_{k_{0}+1})},$
@@
By Lemma 3.5, for sufficiently large $k_{0}$ and any Bowen ball $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ with $z\in\mathbb{F}$, since $B_{M_{k_{0}}}(z,\frac{\epsilon^{\ast}}{2})$ is open, one has
@@ 3.2.1 Upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$
The following lemma allows us to get an upper bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K},d)$.
@@
[Y23, Lemma 3.3,(3)] Let $(X,d,f)$ be a TDS, and let $K\subset M_{f}(X)$ be a non-empty closed subset.
@@
Thus, by Lemma 3.6 we have
@@ 3.2.2 Lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$
The proof for the lower bound of $\overline{\rm mdim}^{B}_{M}(f,G_{K}^{\omega,U},d)$ is similar to that in subsection 3.1.2. So we only give a sketch for the proof.
For any sufficiently small $\epsilon^{\ast}>0$, we show
@@
Let $\mathcal{U}$ be an open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 20\epsilon^{\ast}$, and ${\rm Leb}(\mathcal{U})\geq 5\epsilon^{\ast}$. For $(k,i)\in\mathcal{A}^{\prime}$, by Lemma 3.3 let $\nu_{(k,i)}=\sum_{l=1}^{p_{k,i}}\lambda_{k,i,l}\nu_{l}^{k,i}$ be the finite convex combination with rational coefficients of $f$-ergodic probability measures such that
@@
Repeating the previous argument as in subsection 3.1.2, we choose an integer $n_{(k,i)}^{\prime}$ such that the following conditions are satisfied:
@@
$(2)$ there exist subsets $E_{k,i,l}$, $1\leq l\leq p_{k,i}$, which are $(\lambda_{k,i,l}n_{(k,i)}^{\prime},5\epsilon^{\ast})$-separated satisfying
$E_{k,i,l}\geq\exp\left[\lambda_{k,i,l}n_{(k,i)}^{\prime}\cdot\left(\inf_{\eta\succ\mathcal{U}}h_{\nu_{l}^{k,i}}(f,\eta)-\zeta\right)\right];$
@@
$\mathcal{C}_{(k,i)}\geq\exp\left[n_{(k,i)}(1-\zeta)\cdot(H-2\zeta)\right].$
@@
Set $\mathcal{S}_{(k,0)}=\mathcal{C}_{(k,0)}$, $c_{k,0}=n_{(k,0)}$ for $k\geq 0$. For every $(k,i)\in\mathcal{A}^{\prime}$ and any tuple of points $(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})\in\mathcal{C}_{(k,i)}^{N_{(k,i)}}$, there exists a point $z=z(y_{k,i,1},y_{k,i,2},\ldots,y_{k,i,N_{(k,i)}})$ that $\epsilon_{(k,i)}$-shadows the pieces of orbits
@@
Without loss of generality, we can assume that $\rho_{(k,j)}\to\rho$. Furthermore, one can check that $\rho(\mathbb{F})=1$, and for every sufficiently large $n$ and $B_{n}(z,\frac{\epsilon^{\ast}}{2})\cap\mathbb{F}$, it holds that
@@
Then, by the Lemma 2.1 we have
@@
The more satisfactory results would be to exchange the order of $\limsup_{\epsilon\to 0}$ and $\sup_{\mu\in K}$ (or $\inf_{\mu\in K}$) of Theorems 1.1 and 1.2. However, up to now, the authors remain unaware of whether there exist any counter-examples among infinite entropy systems with the specification property for which the following equalities hold strictly, i.e.,
Inspired by the definition of mean Rényi information dimension [GS21], we define two variants of mean Rényi information dimension of invariant measures.
Let $K$ be a closed set of $M_{f}(X)$. For any $\mu\in K$, we use $M_{K}(\mu)$ to denote the space of sequences of invariant measures within $K$ which converge to $\mu$ in the weak-topology.
Definition 3.1.
For any $\mu\in K$, we define the Bowen and packing Rényi information dimensions of $\mu$ as
$\underline{\rm mrid}_{\mu}^{B}(f,K,d):$ $=\inf_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\liminf_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
$\overline{\rm mrid}_{\mu}^{P}(f,K,d):$ $=\sup_{(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)}\{\limsup_{\epsilon\to 0}\frac{1}{|\log\epsilon|}\inf_{\mathrm{diam}(\xi)<\epsilon}h_{\mu_{\epsilon}}(f,\xi)\},$
respectively.
With the help of Bowen and packing Rényi information dimensions, we establish the variational principles for Bowen and packing metric mean dimensions of saturated sets, whose form are close to the ones of Bowen and packing topological entropies of saturated sets [PS07, ZCC12].
Now we are ready to prove Theorem 1.3. We first prove the variational principle:
$\underline{\rm mdim}_{M}^{B}(f,G_{K},d)=\inf_{\mu\in K}\underline{\rm mrid}_{\mu}^{B}(f,K,d).$
Fix $\mu\in K$. For every $(\mu_{\epsilon})_{\epsilon}\in M_{K}(\mu)$, by Lemma 3.6 it holds that